OpenStudy (anonymous):

In a school of 200 students, 80 students are in the band, 160 students are on sports teams, and 60 students participate in both activities. How many students in the school are neither in band nor on a sports team?

OpenStudy (campbell_st):

number of band only = 80 - 60 number of sport only = 160 - 60 number of both = 60 the number of neither is 200 - the above amounts